Consider the line -3x-7y=8
What is the slope of a line parallel to this line?
What is the slope of a line perpendicular to this line?

Answer by dkppathak(439)

-3x-7y=8
-7y=8+3x
y=8/-7+3/7x
comparing
y=mx+c
m=3/7
when lines are parallel
slop m1= m2
slope of parallel line is 3/7
and slope of perpendicular line is =-7/3